OpenStudy (anonymous):

If log M/N=5 and log P/N=7, what can you say about the relationship between M and P? a. P = 2M b. P = 10M c. P = 100M d. M = 2P

OpenStudy (shamil98):

\[\large \log \frac{ P }{ N } - \log \frac{ M }{ N } = 7 -5\] simplify this first

OpenStudy (shamil98):

the right hand side, what does it then become?

OpenStudy (anonymous):

2?

OpenStudy (shamil98):

yes. now. split the logs on the left hand side using: \[\log \frac{ a }{ b } = \log a - \log b\]

OpenStudy (shamil98):

what do you get?

OpenStudy (anonymous):

i dont know